### v1

@@
250 words in abstract; 3,929 words in main body; 2 figures (0 black and white, 2 colour); 4 tables; 34 references.
@@ 1 Abstract
Materials and Methods: Employing GPT-3.5 we generated and coded 9,606 discharge summaries based on lists of ICD-10 code descriptions of patients with infrequent (generation) codes within the MIMIC-IV dataset. Combined with the baseline training set, this formed an augmented training set. Neural coding models were trained on baseline and augmented data and evaluated on a MIMIC-IV test set. We report micro- and macro-F1 scores on the full codeset, generation codes, and their families. Weak Hierarchical Confusion Matrices were employed to determine within-family and outside-of-family coding errors in the latter codesets. The coding performance of GPT-3.5 was evaluated both on prompt-guided self-generated data and real MIMIC-IV data. Clinical professionals evaluated the clinical acceptability of the generated documents.
Results: Augmentation slightly hinders the overall performance of the models but improves performance for the generation candidate codes and their families, including one unseen in the baseline training data. Augmented models display lower out-of-family error rates. GPT-3.5 can identify ICD-10 codes by the prompted descriptions, but performs poorly on real data. Evaluators note the correctness of generated concepts while suffering in variety, supporting information, and narrative.
Discussion and Conclusion: GPT-3.5 alone is unsuitable for ICD-10 coding. Augmentation positively affects generation code families but mainly benefits codes with existing examples. Augmentation reduces out-of-family errors. Discharge summaries generated by GPT-3.5 state prompted concepts correctly but lack variety, and authenticity in narratives. They are unsuitable for clinical practice.
@@ 2 BACKGROUND AND SIGNIFICANCE
Large-Scale Multi-Labelled Text Classification (LMTC) tasks in Natural Language Processing (NLP) associate input documents with a set of output labels from a large label space, often hierarchically with a big-head long-tail distribution and data sparsity issues.
Medical document coding is the task of assigning structured codes from a medical ontology – e.g., the International Classification of Diseases (ICD) [^1] https://www.who.int/standards/classifications/classification-of-diseases – to clinical documents, a task performed by specially trained hospital staff. Coding consumes human resources that could be allocated to patient care. To ease this burden, research in machine learning and NLP cast medical document coding as an LMTC task [1].
@@
Recently, Large Language Models (LLMs), notably GPT-3.5, have become a new standard for advanced NLP tasks, especially ones reliant on understanding natural language. These models retain and apply background knowledge observed during training, yet can also produce fluent but inaccurate information (known as hallucination) [18]. Writing and coding discharge summaries requires extensive background knowledge making it of interest to explore an LLM’s capability of reading, coding, and generating discharge summaries.
An LLM capable of handling medical text could address the data sparsity issue by synthesising new data. This study aims to investigate the viability of GPT-3.5-generated medical documents for data augmentation in training local neural models and their credibility in clinical settings. We investigate GPT-3.5 given its performance in natural language understanding tasks requiring background knowledge, within an ethical experimental setting for clinical note data (by disabling content monitoring by the service provider). Furthermore, we explore GPT-3.5’s ability to code real discharge summaries and self-generated text.
@@
We queried GPT-3.5 (gpt-3.5-turbo) [^3] https://platform.openai.com/docs/models/gpt-3-5, from hereon referred to as G, through the OpenAI Python API [^4] https://platform.openai.com/docs/api-reference?lang=python to generate a patient discharge summary based on specific conditions and procedures represented by ICD-10-CM and ICD-10-PCS code descriptions from gold standard labels associated with MIMIC-IV discharge summaries (from the table hosp/d_icd_diagnoses.csv.gz). These produced label descriptions chosen to represent realistic combinations of conditions and procedures are not considered a part of the MIMIC-IV dataset. Note that sharing access to MIMIC data via any online API is prohibited [^5] https://physionet.org/news/post/415666Our method was consulted with and approved by PhysioNet, as we merely use the descriptions of attached codes.
While various dataset splits have been proposed since the release of coded discharge summaries in MIMIC-IV, we specifically chose not to follow the one proposed by Edin et al. [24] due to its exclusion of the long tail, which contrasts with our aim of addressing this aspect through generation techniques. Instead, we adopted the dataset split proposed by Nguyen et al. [25] that preserved the long-tail, aligning better with our focus. Nonetheless, we utilised the implementation of common ICD coding models produced by Edin et al. [24] for our analysis.
@@ 2.1 Label Selection
For these 195 codes we produced a list of codes belonging to their families (identified by the head of the code). Families with at least one relatively frequent code (population $>$ 100 in training) and at least one code exclusive to the test set (zero-shot) were retained leaving 16 families. We randomly selected 10 of these to generate from. For each family, we gathered its descendant codes appearing in the training set (139) and the zero-shot codes (16) totalling 155 codes. Of these, 114 codes had a population lower than 100 and are henceforth termed generation codes (list of codes available in Supplementary Material).
@@ 2.2 Preparation of Samples for Generation
In the Nguyen’s MIMIC-IV training set, we found documents with the 98 relevant few-shot codes (the 16 zero-shot were by-definition absent). Some documents contained multiple relevant codes. We have collected documents for each of the relevant codes and cloned them to bring their population up to 100. To reduce duplication, we have randomly dropped up to 5 of the assigned non-relevant labels within clones to create the new set of labels for generation (referred to as the silver standard).
We identified documents containing the siblings of the 16 zero-shot labels. A silver standard set was created for each of these documents substituting the sibling code with the zero-shot code (similar to the zero-shot approach in [10]). If multiple siblings were present, a random one was replaced with the zero-shot code.
This resulted in 9,606 input sets of labels – 6,779 unique and 2,827 duplicated.
@@ 2.3 Generation
Natural Language Generation is the task of producing natural language text based on a set of input data. We used the model “gpt-3.5-turbo-0613” given its wide recognition in the field of LLMs, relative cost-effectiveness, and time efficiency (compared to gpt-4). For distinct label sets, we utilised a temperature (parameter in the 0-1 range controlling randomness) of 0 to produce deterministic outputs. A temperature of 0.1 was employed for duplicates, allowing output variation.
Within the prompt (available in Supplementary Material), we state the task of writing a discharge summary for a patient with a list of standard descriptions of their conditions and procedures based on our silver standard. We also add further specifications:
@@
  • Anonymisation was required for personal and location data (due to uncertainty of the anonymity of the data used in training of GPT-3.5), maintaining numeric information when relevant despite potential removal in preprocessing;
@@
The generated documents were processed to find the discharge diagnoses and extract the assigned codes with a regular expression.
In total, we have generated 9,606 synthetic training documents. We removed all mentions of ICD-10 codes and pre-processed the documents the same as the baseline data. Then we merged these documents with the baseline training set (110,442 MIMIC-IV documents), forming the augmented training set (120,048 documents). The baseline and augmented settings used the same validation and test sets with 4,017 and 7,851 real documents, respectively.
Analysis between the generated text, and MIMIC-IV ICD-10-coded discharge summaries (the entire dataset, and the subset used for source label sets used in generation) showed differences in word count (per label and overall – Figures 1a and 1b) between synthetic and real data while retaining similar distributions of label counts (Figure 1c). The synthetic discharge summaries tend to be shorter which could be the result of GPT-3.5’s 4,096 token limit. Upon publication, we intend to make the synthetic discharge available.
Figure 1: A comparison between the discharge summary data in MIMIC-IV, the source data, and the generated discharge summaries. Subfigures 1a and 1b focus on the number of words in documents, indicating that the GPT-generated data generally contains fewer words overall and per assigned label compared to the real data from MIMIC-IV. Subfigure 1c demonstrates that, although there’s a variance in document size, the distribution of the number of labels per document remains relatively similar across the datasets.
@@ 2.5.1 Local Neural Model Evaluation
The models’ test performance was evaluated using standard information retrieval metrics commonly used in LMTC tasks – micro- and macro-averaged Precision, Recall, and F1 scores
Micro-averaging assigns equal weight to each prediction, favouring high-population classes (e.g., hypertension). Macro-averaging, in contrast, computes the performance for each unique label and averages across the label space, giving each label’s average result equal weight regardless of their population. This highlights poor performance in less common classes. Our primary evaluation metrics common with the majority of previous work are micro-F1 and macro-F1 scores. Metrics are further explained in Supplementary Material.
@@ 2.5.2 GPT’s coding on real clinical notes
For reproducibility, we specified the model version and the API version as “gpt-3.5-turbo-0613” and “2023-03-15-preview”, respectively. All parameters were set to zero for deterministic responses from G, including temperature, top P, frequency penalty, and presence penalty. The system prompt directed G to act as a clinician assigning ICD-10 diagnosis codes to clinical notes, specifying the expected output format as JSON objects with keys “diagnosis” and “icd_code”. Table 1 presents all hyperparameter details. Our code implementation can be found in our Github repository  [^8] https://github.com/EdinburghClinicalNLP/chatgpt_icd_coding.
Parameter
Value
Model Name
gpt-3.5-turbo-0613
API Version
2023-03-15-preview
Temperature
0
Top P
0
Frequency Penalty
0
Presence Penalty
0
System Prompt
You are a clinician and your task is to assign diagnoses and the associated ICD-10 clinical codes to a given clinical note. You should output an array of JSON objects containing the diagnosis name and the ICD-10 code, structured as diagnosis and icd_code.
Table 1: Azure API call hyperparameters.
@@
In our evaluation of GPT’s performance, we have also employed hierarchical evaluation techniques – set-based hierarchical evaluation  [31] and Count-Preserving Hierarchical Evaluation (CoPHE) [32]. These metrics award partial credit to mispredicted labels by extending prediction and gold standard sets with their ancestor labels. While set-based evaluation ancestor labels track only presence of descendants, in CoPHE, ancestor labels link to the count of descendants, penalising over- and under-predictions within code families.
Comparing set-based and CoPHE results helps to evaluate the model’s tendency to over-/under-predict. A lower CoPHE score indicates this phenomenon. See Supplementary Materials for details on calculating hierarchical scores.
@@ 2.5.4 Acceptability of Generated Data in Clinical Practice
Additionally, they separately evaluated correctness and informativeness for both non-low-resource and low-resource labels to gauge G’s ability to generate low-resource data. Scores from 1 (failure to perform) to 5 (perfect performance) were assigned to each metric, with accompanying comments justifying the score.
@@ 3.1 Local Neural Model Evaluation
We assessed performance across three code subsets: the entire codeset in MIMIC-IV ($overall$), a restricted set encompassing our candidate generation labels and their related codes ($f$), and a further subset ($f(gen)$) containing only the 114 low-population candidate generation labels. Results are shown in Table 2. Baseline results for CAML and LAAT align closely with prior findings by Nguyen et al. [25] for $overall$ metrics. While Nguyen et al. [25] did not report on Multi-Res CNN, its performance trends were similar to CAML and LAAT in comparison to Edin et al. [24] on non-filtered codesets. LAAT excels in micro-F1, Multi-Res CNN leads in macro-F1, and CAML generally lags behind. Baseline models outperform augmented ones in $overall$ micro-F1, a common observation when enhancing lower-resource label performance. Nonetheless, macro-F1 scores improved for two out of three models within the $overall$ codeset and for all models on $f$ and $f(gen)$. Multi-Res CNN and CAML macro-F1 scores display sizable relative improvement (26% and 78% respectively) in $f(gen)$.
Augmented models performed on par with or outperformed baseline models in micro-F1 scores for $f$ and $f(gen)$. Augmented Multi-Res CNN outperforms its baseline in micro-F1 for both $f$ and $f(gen)$, indicating benefits for the code family from augmenting less-populous members. Augmented LAAT shows improvement in macro-F1 in both $f$ and $f(gen)$ but lags in micro-F1
Comparing within-family and out-of-family errors (Table 3), augmented models generally exhibit fewer out-of-family errors on $f$ and $f(gen)$. An exception to this is augmented Multi-Res CNN in $f(gen)$, whose slight increase in OOF came with a sizable reduction in within-family error.
@@
Experiment $mi_{overall}$ $ma_{overall}$ $mi_{f}$ $ma_{f}$ $mi_{f(gen)}$ $ma_{f(gen)}$
CAML baseline 53.651 3.866 38.431 3.027 17.413 6.635
CAML augmented 53.544 3.899 38.410 3.780 20.680 11.858
LAAT baseline 57.285 6.181 43.586 4.959 26.790 14.478
LAAT augmented 57.176 6.091 43.364 5.381 25.696 14.978
Multi-Res CNN baseline 55.663 6.402 40.159 5.039 26.801 13.923
Multi-Res CNN augmented 54.694 6.464 42.688 5.846 30.390 17.681
Table 2: A comparison between local neural network models trained on baseline and augmented training sets is evaluated using micro- and macro-averaged $F1$ scores ($mi$ and $ma$ respectively) on three codesets – $overall$ on all codes present in MIMIC-IV; $f$ comprising all codes within the families we chose for generation; and $f(gen)$ corresponding to candidate codes used in generation with a population of at most 100 in the training set. The highest score in each metric for each model pair (baseline versus augmented) is highlighted in bold.
[…]
CAML baseline 66.525 25.045 83.809 9.831
CAML augmented 65.983 23.768 79.785 9.169
LAAT baseline 58.568 28.351 74.033 12.197
LAAT augmented 55.932 29.783 73.650 11.981
Multi-Res CNN baseline 52.720 32.406 69.676 15.242
Multi-Res CNN augmented 49.653 32.359 70.413 10.216
Table 3: Weak Hierarchical Confusion Matrix (WHCM) error rates of local neural network models trained on the original (baseline) training set versus on the training set augmented with synthetic data (augmented). Performance on the common test set is reported using the macro-averaged proportion of errors that were Out-of-Family (OOF) and within-family (In-Family). Errors are analysed for two codesets – $f$ on the codes within the families of codes used for generation; and $f(gen)$ corresponding to codes in $f$ with a population of at most 100 in the baseline training set. The best (lowest) error rate for each error type for each model pair (baseline versus augmented) is presented in bold.

### v2

@@
250 words in the abstract; 4,152 words in the main body; 4,428 words in the main body + Funding, Competing Interests, and Author Contribution sections; 4 figures (1 black and white, 3 colour); 4 tables; 34 references.
@@ 1 Abstract
Materials and Methods: Employing GPT-3.5 we generated and coded 9,606 discharge summaries based on lists of ICD-10 code descriptions of patients with infrequent (or generation) codes within the MIMIC-IV dataset. Combined with the baseline training set, this formed an augmented training set. Neural coding models were trained on baseline and augmented data and evaluated on a MIMIC-IV test set. We report micro- and macro-F1 scores on the full codeset, generation codes, and their families. Weak Hierarchical Confusion Matrices determined within-family and outside-of-family coding errors in the latter codesets. The coding performance of GPT-3.5 was evaluated on prompt-guided self-generated data and real MIMIC-IV data. Clinicians evaluated the clinical acceptability of the generated documents.
Results: Data augmentation results in slightly lower overall model performance but improves performance for the generation candidate codes and their families, including one absent from the baseline training data. Augmented models display lower out-of-family error rates. GPT-3.5 identifies ICD-10 codes by their prompted descriptions, but underperforms on real data. Evaluators highlight the correctness of generated concepts while suffering in variety, supporting information, and narrative.
Discussion and Conclusion: While GPT-3.5 alone given our prompt setting is unsuitable for ICD-10 coding, it supports data augmentation for training neural models. Augmentation positively affects generation code families but mainly benefits codes with existing examples. Augmentation reduces out-of-family errors. Documents generated by GPT-3.5 state prompted concepts correctly but lack variety, and authenticity in narratives.
@@ 2 BACKGROUND AND SIGNIFICANCE
Large-Scale Multi-Labelled Text Classification (LMTC) tasks in Natural Language Processing (NLP) associate input documents with a set of output labels from a large label space, often hierarchically with a big-head long-tail distribution and data sparsity issues. Medical document coding is the task of assigning structured codes from a medical ontology – e.g., the International Classification of Diseases (ICD) [^1] https://www.who.int/standards/classifications/classification-of-diseases – to clinical documents, a task performed by specially trained hospital staff. Coding consumes human resources that could be allocated to patient care. To ease this burden, research in machine learning and NLP cast medical document coding as an LMTC task [1].
@@
Recently, Large Language Models (LLMs), notably GPT-3.5, have become a new standard for advanced NLP tasks, especially ones reliant on understanding natural language. These models retain and apply background knowledge observed during training, yet can also produce fluent but inaccurate information (known as hallucination) [18]. Writing and coding discharge summaries require extensive background knowledge making it of interest to explore an LLM’s capability of reading, coding, and generating discharge summaries. An LLM capable of handling medical text could address the data sparsity issue by synthesising new data. This study aims to investigate the viability of GPT-3.5-generated medical documents for data augmentation in training local neural models and their credibility in clinical settings. We investigate GPT-3.5 given its performance in natural language understanding tasks requiring background knowledge, within an ethical experimental setting for clinical note data (by disabling content monitoring by the service provider). Furthermore, we explore GPT-3.5’s ability to code real discharge summaries and self-generated text.
@@
We queried GPT-3.5 (gpt-3.5-turbo) [^3] https://platform.openai.com/docs/models/gpt-3-5, from hereon referred to as G, through the OpenAI Python API [^4] https://platform.openai.com/docs/api-reference?lang=python to generate patient discharge summaries based on specific conditions and procedures represented by ICD-10-CM and ICD-10-PCS code descriptions from gold standard labels associated with MIMIC-IV discharge summaries (from the table hosp/d_icd_diagnoses.csv.gz). These label combinations were chosen to closely follow real scenarios to correspond to correlations between real labels. Note that sharing MIMIC data via any online API is prohibited [^5] https://physionet.org/news/post/415666Our method was consulted with and approved by PhysioNet, as we merely use the descriptions of attached codes, which are not considered part of the dataset..
While various dataset splits have been proposed since the release of coded discharge summaries in MIMIC-IV, we specifically chose not to follow the one proposed by Edin et al. [24] due to its exclusion of the long tail, which contrasts with our aim of addressing this aspect through generation techniques. Instead, we adopted the dataset split proposed by Nguyen et al. [25] (Table 1)that preserved the long-tail, aligning better with our focus. Nonetheless, we utilised the implementation of common ICD coding models produced by Edin et al. [24] for our analysis.
Table 1: MIMIC-IV Nguyen dataset split. Zero-Shot corresponds to labels absent from the training set. Few-Shot corresponds to labels appearing at least once but no more than five times in the training set.
Train Dev Test
# Documents 110,442 4,017 7,851
# Labels (Total) 1,784,304 65,516 124,518
# Labels (Unique) 25,230 6,738 9,159
# Labels (Unique Zero-Shot) N/A 291 587
# Labels (Unique Few-Shot) 15,300 919 1,646
@@ 2.1 Label Selection
For these 195 codes we produced a list of codes belonging to their families (identified by the head of the code). Families with at least one relatively frequent code (population $>$ 100 in training) and at least one code exclusive to the test set (zero-shot) were retained leaving 16 families. The constraint of having at least one frequent code was included to explore the scenario where the model may predict a label due to its dominance in the population, and the increase in population of other labels potentially leading to more confusion.
From these 16 families we randomly selected 10 to generate from (E10, G43, H35, H81, S00, S02, S06, T82, T84, T85). Of these, 114 codes had a population lower than 100 and are henceforth termed generation codes (a list of codes is available in Supplementary Material 1).
@@ 2.2 Preparation of Samples for Generation
As there exist co-relations among labels (e.g., different complications of type 1 diabetes, cancer co-relating with the presence of chemotherapy), rather than creating random combinations of ICD codes we opted to work our way back from existing real scenarios.
In Nguyen’s MIMIC-IV training set, we found documents with the 98 relevant few-shot codes (the 16 zero-shot were by-definition absent). Some documents contained multiple relevant codes. We have collected documents for each of the relevant codes and cloned them to bring their population up to 100. To increase variety, we have randomly dropped up to 5 of the assigned non-relevant labels within clones to create the new set of labels for generation (referred to as the silver standard).
We identified documents containing the siblings of the 16 zero-shot labels. A silver standard set was created for each of these documents substituting the sibling code with the zero-shot code (similar to the zero-shot approach in [10]). If multiple siblings were present, a random one was replaced with the zero-shot code. This resulted in 9,606 input sets of labels – 6,779 unique and 2,827 duplicated.
@@ 2.3 Generation
Natural Language Generation is the task of producing natural language text based on a set of input data. We used the model “gpt-3.5-turbo-0613” given its wide recognition in the field of LLMs, relative cost-effectiveness, and time efficiency (compared to gpt-4). We utilised a temperature (parameter in the 0-1 range controlling randomness) of 0 to produce deterministic outputs. We set the temperature to 0.1 for duplicates, allowing output variation.
Within the prompt (see Supplementary Material 2), we specified to write a discharge summary for a patient with a list of standard descriptions of their conditions and procedures based on our silver standard. We added further specifications:
Figure 1: An example generation of a synthetic discharge via GPT-3.5
@@
  • Anonymisation was required for personal and location data (due to uncertainty of the anonymity of the data used in training of GPT-3.5), maintaining numeric information when relevant despite potential removal in pre-processing;
@@
An example of the generation process can be seen in Figure 1.
The generated documents were processed to find the discharge diagnoses and extract the assigned codes with a regular expression. In total, we have generated 9,606 synthetic training documents. We removed all mentions of ICD-10 codes and pre-processed the documents the same as the baseline data. Then we merged these documents with the baseline training set (110,442 MIMIC-IV documents), forming the augmented training set (120,048 documents). The baseline and augmented settings used the same validation and test sets with 4,017 and 7,851 real documents, respectively.
Analysis between the generated text, and MIMIC-IV ICD-10-coded discharge summaries (the entire dataset, and the subset used for source label sets used in generation) showed differences in word count (per label and overall – Figures 2a and 2b) between synthetic and real data while retaining similar distributions of label counts (Figure 2c). The synthetic discharge summaries tend to be shorter which could be the result of GPT-3.5’s 4,096 token limit. The synthetic discharge summaries have been accepted for publication and will be and will be made available by PhysioNet.
Figure 2: A comparison between the discharge summary data in MIMIC-IV, seed MIMIC discharge summaries for generation (the source data), and the generated discharge summaries. Subfigures 2a and 2b focus on the number of words in documents, indicating that the GPT-generated data generally contains fewer words overall and per assigned label compared to the real data from MIMIC-IV. Subfigure 2c demonstrates that, although there’s a variance in document size, the distribution of the number of labels per document remains relatively similar across the datasets.
@@ 2.5.1 Local Neural Model Evaluation
The models’ test performance was evaluated using standard information retrieval metrics commonly used in LMTC tasks – micro- and macro-averaged Precision, Recall, and F1 scores.
Micro-averaging assigns equal weight to each prediction, favouring high-population classes (e.g., hypertension). Macro-averaging, in contrast, computes the performance for each unique label and averages across the label space, giving each label’s average result equal weight regardless of their population. This highlights poor performance in less common classes. Our primary evaluation metrics common with the majority of previous work are micro-F1 and macro-F1 scores. Metrics are further explained in Supplementary Material 3.
@@ 2.5.2 GPT’s coding on real clinical notes
For reproducibility, we specified the model version and the API version as “gpt-3.5-turbo-0613” and “2023-03-15-preview”, respectively. All parameters were set to zero for deterministic responses from G, including temperature, top P, frequency penalty, and presence penalty. The system prompt directed G to act as a clinician assigning ICD-10 diagnosis codes to clinical notes, specifying the expected output format as JSON objects with keys “diagnosis” and “icd_code”. Refer to Figure 3 for all hyperparameter details. Our code implementation can be found in our Github repository  [^8] https://github.com/EdinburghClinicalNLP/chatgpt_icd_coding.
@@
In our evaluation of GPT’s performance, we have also employed hierarchical evaluation techniques – set-based hierarchical evaluation  [31] and Count-Preserving Hierarchical Evaluation (CoPHE) [32]. These metrics award partial credit to mispredicted labels by extending prediction and gold standard sets with their ancestor labels. While set-based evaluation ancestor labels track only the presence of descendants, in CoPHE, ancestor labels link to the count of descendants, penalising over- and under-predictions within code families.
Comparing set-based and CoPHE results helps to evaluate the model’s tendency to over-/under-predict. A lower CoPHE score indicates this phenomenon. See Supplementary Materials 4 for details on calculating hierarchical scores.
@@ 2.5.4 Acceptability of Generated Data in Clinical Practice
Additionally, they separately evaluated correctness and informativeness for both non-low-resource and low-resource labels to gauge G’s ability to generate low-resource data. Scores from 1 (failure to perform) to 5 (perfect performance) were assigned to each metric, with accompanying comments justifying the score. An example evaluation by a clinician can be seen in Figure 4.
Figure 4: An example evaluation of a synthetic discharge summary by a clinical expert.
@@ 3.1 Local Neural Model Evaluation
We assessed performance across three code subsets: the entire codeset in MIMIC-IV ($overall$), a restricted generation set ($f_{(\text{gen})}$) containing only the 114 low-population candidate generation labels, and a set of codes from the code families present in $f_{(\text{gen})}$ ($f$). Results are shown in Table 2. Baseline results for CAML and LAAT align closely with prior findings by Nguyen et al. [25] for $overall$ metrics. While Nguyen et al. [25] did not report on Multi-Res CNN, its performance trends were similar to CAML and LAAT in comparison to Edin et al. [24] on non-filtered codesets. LAAT excels in micro-F1, Multi-Res CNN leads in macro-F1, and CAML generally lags behind. Baseline models outperform augmented ones in $overall$ micro-F1, a common observation when enhancing lower-resource label performance. Nonetheless, macro-F1 scores improved for two out of three models within the $overall$ codeset and for all models on $f$ and $f_{(\text{gen})}$. Multi-Res CNN and CAML macro-F1 scores display sizable relative improvement (26% and 78% respectively) in $f_{(\text{gen})}$.
Augmented models performed on par with or outperformed baseline models in micro-F1 scores for $f$ and $f_{(\text{gen})}$. Augmented Multi-Res CNN outperforms its baseline in micro-F1 for both $f$ and $f_{(\text{gen})}$, indicating benefits for the code family from augmenting less-populous members. Augmented LAAT shows improvement in macro-F1 in both $f$ and $f_{(\text{gen})}$ but lags in micro-F1. LAAT’s performance may have been biased towards high-population classes and the augmentation’s boosting of low-frequency classes (misrepresenting their frequency) may have introduced confusion. Apart from having a recurrent encoder (Bi-LSTM), the LAAT model employed in this experiment is about twice the size of the Multi-Res CNN (21.9M versus 11.9M parameters). This added model complexity may have enabled better performance on already frequent labels, but increased the need for more examples of lower-resource labels.
Comparing within-family and out-of-family errors (Table 2), augmented models generally exhibit fewer out-of-family errors on $f$ and $f_{(\text{gen})}$. An exception to this is augmented Multi-Res CNN in $f_{(\text{gen})}$, whose slight increase in OOF came with a sizable reduction in within-family error.
@@
Table 2: A comparison between local neural network models (MRCNN stands for Multi-Res CNN) trained on baseline ($base$) and augmented ($aug$) training sets is evaluated using micro- and macro-averaged $F1$ scores ($mi$ and $ma$ respectively) on three codesets – $overall$ on all codes present in MIMIC-IV; $f$ comprising all codes within the families we chose for generation; and $f_{(\text{gen})}$ corresponding to candidate codes used in generation with a population of at most 100 in the training set. The highest score in each metric for each model pair (baseline versus augmented) is highlighted in bold. Weak Hierarchical Confusion Matrix (WHCM) error rates are produced for codesets $f$ and $f_{(\text{gen})}$. Performance on the common test set is reported using the macro-averaged proportion of errors that were Out-of-Family (OOF) and within-family (IF). The best (lowest) error rate for each error type for each model pair (baseline versus augmented) is presented in bold.
Experiment F1 $\uparrow$ WHCM error $\downarrow$
$mi_{overall}$ $ma_{overall}$ $mi_{f}$ $ma_{f}$ $mi_{f(\text{gen})}$ $ma_{f(\text{gen})}$ $OOF_{f}$ $IF_{f}$ $OOF_{f(\text{gen})}$ $IF_{f(\text{gen})}$
CAMLbase 53.65 3.87 38.43 3.03 17.41 6.64 66.53 25.05 83.81 9.83
CAMLaug 53.54 3.90 38.41 3.78 20.68 11.86 65.98 23.77 79.79 9.17
LAATbase 57.29 6.18 43.59 4.96 26.79 14.48 58.57 28.35 74.03 12.20
LAATaug 57.18 6.09 43.36 5.38 25.70 14.98 55.93 29.78 73.65 11.98
MRCNNbase 55.66 6.40 40.16 5.04 26.80 13.92 52.72 32.41 69.68 15.24
MRCNNaug 54.69 6.46 42.69 5.85 30.39 17.68 49.65 32.36 70.41 10.22